OpenStudy (anonymous):

Find dy/dx y=xlog2(x)

OpenStudy (ybarrap):

Do you know the product rule? You have that y is a PRODUCT of two functions of x. The first function is just "x" the second is log(x).

OpenStudy (anonymous):

I got here: lny=lnx + ln log2(x)

OpenStudy (ybarrap):

You are using log with a base of 2, right?

OpenStudy (anonymous):

yes

OpenStudy (ybarrap):

Ok, so rewrite that as $$ \log_2(x) = {\ln(x)\over \ln(2)} $$

OpenStudy (ybarrap):

Where ln(x) is the natural log of x

OpenStudy (ybarrap):

$$ \cfrac{d}{dx}x\log_2(x)=\cfrac{d}{dx}x{\ln(x)\over \ln(2)}=\cfrac{1}{\ln(2)}\cfrac{d}{dx}\left [x\ln(x)\right ] $$ Now apply product rule

OpenStudy (ybarrap):

do you know the product rule?

OpenStudy (anonymous):

no

OpenStudy (ybarrap):

You have y that is a product of two functions of x so you need to apply the product rule $$ y=f_1(x)\times f_2(x)\\ \cfrac{d}{dx}y=f_2(x)\cfrac {d}{dx}f_1(x)+f_1(x)\cfrac {d}{dx}f_2(x)\\ $$ Where $$ f_1(x)=x $$ and $$ f_2(x)=\log_2(x) $$ So now you just need to find $$ \cfrac{d}{dx}f_1(x) $$ and $$ \cfrac{d}{dx}f_2(x) $$ Does that all make sense?

OpenStudy (ybarrap):

In words, the derivatives of y is log2(x) times the derivative of "x" PLUS "x" times the derivative of log2(x). That is the product rule.

OpenStudy (triciaal):